OpenStudy (anonymous):

The smaller of two numbers is 12 less than three times the larger. If their difference is 2, what is the larger number?

OpenStudy (anonymous):

let x be the smaller number let y be the larger number then 3y -12 = x (1) y -2 = x (2) sub (1) into (2) 3y-12 =y -2 2y = 10 y = 5 sub y= 5 into (2) (5) -2 = x x = 3 thus the numbers are 5 and 3